If the interest rate on euro-denominated assets is 13 percent and it is 15 percent on peso-denominated assets


Question 18 

If the interest rate on euro-denominated assets is 13 percent and it is 15 percent on peso-denominated assets, and if the euro is expected to appreciate at a 4 percent rate against peso, for Manuel the Mexican the expected rate of return on euro-denominated assets is _______% 


Question 19

According to the purchasing power parity theory, a rise in the United States price level of 5 percent, and a rise in the Mexican price level of 6 percent cause the dollar to appreciate by  _______% (put a negative sign if it depreciates) relative to the peso.
Answer #1

Question 18

Assuming the rate of exchange as 1 Euro = 25.92 Mexican Peso ( as on date July 29th ' 2020)

Interest Rate in Euro denominated assets : 13%

Interest Rate in Peso denominated assets : 15%

Now taking an example :

If Manuel from Mexico has 1000 Peso, he may have two investment alternatives:

1. 1000 Peso invested in Peso denominated assets with return of 150 Peso (15%)

2. (1000/25.92) Euros [at the exchange rate given above] which equals 38.58 Euros. This will give him 13% interest which amounts to 5.015 Euros. Applying the original exchange rate, we may get a return of 130 Peso (5.015 x 25.92) {13% return}.

Now if the value of Euro appreciates against Peso, then the revised exchange rate would be calculated as:-

1 Euro= (25.92 + 4% of 25.92)Peso

1 Euro= 26.95 Peso

Applying the revised exchange rate on interest amount in Euros of 5.015 Euros, we get 135.10 Peso which gives Manuel the Mexican an effective return of 13.51% on Euro-denominated assets.

Question 19 As per the given information,

Rise in United States price level is 5%

Rise in Mexican price level is 6%

According to Purchasing Power Parity rate between United States and Mexico as per 2019 is 1 USD = 9.309 Mexican Peso.

After the rise in the price level in United States and Mexico:

USD 1 * 1.05 = 9.309 * 1.06 Mexican Peso

1.05 USD = 9.87 Mexican Peso ( After a rise in price level of 5% in US and 6% in Mexico)

Therefore, 1 USD= 9.87 / 1.05

= 9.397 Mexican Peso

Hence, the change in value of USD relative to Peso is

or 0.945%.

The USD appreciates by 0.945% relative to Peso.